# Utility-Learning Tension in Self-Modifying Agents

### v1

@@ Abstract
As systems trend toward superintelligence, a natural modeling premise is that agents can self-improve along every facet of their own design. We formalize this with a five-axis decomposition and a decision layer, separating incentives from learning behavior and analyzing axes in isolation. Our central result identifies and introduces a sharp utility–learning tension, the structural conflict in self-modifying systems whereby utility-driven changes that improve immediate or expected performance can also erode the statistical preconditions for reliable learning and generalization. Our findings show that distribution-free guarantees are preserved iff the policy-reachable model family is uniformly capacity-bounded; when capacity can grow without limit, utility-rational self-changes can render learnable tasks unlearnable. Under standard assumptions common in practice, these axes reduce to the same capacity criterion, yielding a single boundary for safe self-modification. Numerical experiments across several axes validate the theory by comparing destructive utility policies against our proposed two-gate policies that preserve learnability.
@@ 1 Introduction
Evidence for this shift already exists. Reinforcement learning and meta-learning instantiate constrained self-change (Sutton and Barto,, 2018; Finn et al.,, 2017; Rajeswaran et al.,, 2019; Hospedales et al.,, 2022), while open-ended pipelines iterate code edits and tools (Zhang et al.,, 2025). Decision-theoretic proposals investigate provably utility-improving modifications (Schmidhuber,, 2005), safety analyses document pathologies (Orseau and Ring,, 2011), and metagoal frameworks aim to stabilize goal evolution (Goertzel,, 2024). What remains underdeveloped is a learning-theoretic account of post-modification behavior: when do seemingly rational self-changes preserve the conditions under which learning is possible, and when do they destroy them? We prove a policy-level learnability boundary: distribution-free PAC guarantees are preserved iff the policy-reachable family has uniformly bounded capacity. A simple Two-Gate guardrail (validation margin $\tau$ + capacity cap $K(m)$) keeps trajectories on the safe side and yields a VC-rate oracle inequality.
@@
  • Policy boundary (iff). Under standard i.i.d. assumptions, distribution-free PAC learnability is preserved under self-modification iff the policy-reachable family has uniformly bounded capacity (VC/pseudodim).
  • Axis reductions. Architectural and metacognitive edits reduce to induced hypothesis families; substrate changes matter only via the induced family. Hence the boundary depends solely on the supremum capacity of the reachable family.
  • Two-Gate guardrail. A computable accept/reject rule (validation improvement by margin $\tau$ + capacity cap $K(m)$) ensures monotone true-risk steps and an oracle inequality at VC rates for the final predictor.
@@ 2.1 Decision-theoretic self-modification and safety
Gödel Machines give a proof-based framework for globally optimal self-modification under a utility function (Schmidhuber,, 2005). Safety analyses document pathologies for self-modifying agents, including reward hacking and self-termination (Orseau and Ring,, 2011). Open-ended empirical systems iterate code and toolchain edits with benchmark gains but without proof obligations (Zhang et al.,, 2025). Proposals for metagoals aim to stabilize or moderate goal evolution during self-change (Goertzel,, 2024). While these frameworks establish decision-theoretic foundations, they do not provide learning-theoretic guarantees about post-modification generalization. We study the learning-theoretic state of the agent after such edits.
@@ 2.2 Modern mechanisms for self-improvement
Contemporary machine learning exhibits constrained forms of self-modification across multiple dimensions. Neural architecture search explores architectural topologies through differentiable, evolutionary, and reinforcement approaches (Liu et al.,, 2019; Elsken et al.,, 2019; Zoph and Le,, 2017; Real et al.,, 2019). Automated machine learning systems perform pipeline and hyperparameter search and can trigger optimizer and model-family switches (Hutter et al.,, 2019; Feurer and Hutter,, 2019; Li et al.,, 2017). Population-based training simultaneously evolves hyperparameters and weights across a population of models (Jaderberg et al.,, 2017).
Meta-learning adapts optimizers, initializations, and inductive biases across tasks (Finn et al.,, 2017; Rajeswaran et al.,, 2019; Hospedales et al.,, 2022). Reinforcement learning and multi-armed bandits provide policies for selecting modifications and exploration strategies (Sutton and Barto,, 2018; Auer et al.,, 2002; Lai and Robbins,, 1985; Slivkins,, 2019). Representation growth through mixture of experts and adapters, and the use of external memory and retrieval, expand the effective function family and computation available at inference (Fedus et al.,, 2022; Houlsby et al.,, 2019; Hu et al.,, 2021; Graves et al.,, 2014, 2016; Lewis et al.,, 2020; Schick et al.,, 2023). Continual learning addresses sequential task acquisition while mitigating catastrophic forgetting (Kirkpatrick et al.,, 2017; Parisi et al.,, 2019; Van de Ven and Tolias,, 2022).
These mechanisms instantiate partial self-modification—adapting specific components while keeping the learning framework itself fixed. In contrast, true self-modifying agents can rewrite any axis of their design. In our framework, these mechanisms traverse representational, architectural, algorithmic, and metacognitive axes, and we establish when such traversals preserve or destroy learnability.
@@ 2.3 Learning theory for adaptive systems
PAC learning provides distribution-free guarantees under a fixed hypothesis class and algorithm (Shalev-Shwartz and Ben-David,, 2014; Mohri et al.,, 2018; Blumer et al.,, 1989; Vapnik,, 1998; Hanneke et al.,, 2024). Online learning theory establishes regret bounds for adaptive algorithms (Shalev-Shwartz,, 2012; Hazan,, 2016; Cesa-Bianchi and Lugosi,, 2006), but assumes the learning mechanism remains fixed. Transformation-invariant learners extend instance equivalence while keeping the learning mechanism fixed (Shao et al.,, 2022). Predictive PAC relaxes data assumptions with a fixed learner (Pestov,, 2010), and iterative improvement within constrained design spaces admits PAC-style analysis (Attias et al.,, 2025).
Information-theoretic approaches bound generalization for adaptive and meta-learners via mutual information (Jose and Simeone,, 2021; Chen et al.,, 2021; Wen et al.,, 2025). Stability connects optimization choices to generalization bounds (Bousquet and Elisseeff,, 2002; Hardt et al.,, 2016). All of these results assume architectural invariance: the hypothesis class, update rule, or computational model is fixed ex ante. We remove this assumption and characterize when self-modification preserves PAC learnability.
@@ 2.4 Computability and the substrate
Church–Turing-equivalent substrates preserve solvability up to simulation overhead, whereas strictly weaker substrates with finite memory can forfeit learnability of classes that are otherwise PAC-learnable; stronger-than-Turing models change the problem class under discussion (Akbari and Harrison-Trainor,, 2024). This motivates treating substrate edits separately from architectural or representational changes and clarifies when invariance should be expected.
@@ 3 Setup and Five-Axis Decomposition
$\ell_{t}=(A_{t},H_{t},Z_{t},F_{t},M_{t})\in\underbrace{\mathcal{L}}_{\mathcal{A}\times\mathcal{H}\times\mathcal{Z}\times\mathcal{F}\times\mathcal{M}},$
where $A$ (algorithmic), $H$ (representational), $Z$ (architectural), $F$ (substrate), and $M$ (metacognitive) are the five axes.
A possibly stochastic modification map $\Phi:\mathcal{L}\times\mathcal{D}\!\to\!\mathcal{L}$ updates the system via $\ell_{t+1}=\Phi(\ell_{t},D_{t})$. For $X\!\in\!\{A,H,Z,F,M\}$ with state space $\mathcal{X}$,
$X_{t+1}=\Phi_{X}(X_{t},D_{t},\theta_{X,t}),\qquad\theta_{X,t}\in\Theta_{X}.$
A reasonable utility $u$ is computable from finite state and data (for example empirical or validation risk with optional resource or complexity terms) and normalized to $[0,1]$. A state (and its components) is policy-reachable under $u$ if it can be obtained by iterating $\Phi$ from $\ell_{0}$ using the following decision rule.
Decision rule.
A candidate modification at time $t$ is executed iff there is a formal proof in the agent’s current calculus that it yields an immediate utility increase:
$u\!\big(\Phi(\ell_{t},D_{t}),\mathrm{Env}_{t}\big)\;>\;u\!\big(\ell_{t},\mathrm{Env}_{t}\big).$
We make no assumptions about proof-search efficiency; our results concern the post-modification learner.
Policy-reachable families (general form).
For any axis $X\in\{A,H,Z,F,M\}$, let $\mathcal{X}_{\mathrm{reach}}(u)$ be the set of $X$-states appearing in policy-reachable trajectories under $u$. In particular,
$\mathcal{X}_{\mathrm{reach}}(u)=\big\{\,$ $X^{\prime}:\ \exists\,t\text{ along a policy-reachable }$
$\text{trajectory under }u\text{ and }X_{t}=X^{\prime}\,\big\}.$
@@
Losses are bounded in $[0,1]$. Our general statements hold for any uniform capacity notion that yields distribution-free uniform convergence (e.g., VC-subgraph or pseudodimension). For concreteness we instantiate to the 011 loss, where capacity reduces to the VC dimension VC()($\cdot$). (All bounds remain valid if VC is replaced by $\mathrm{Pdim}$ or VC-subgraph with the usual constants.)
@@
[…]
@@
Throughout, we analyze one axis at a time while holding others fixed. Under Church–Turing-equivalent substrates $F$, learnability refers to classical (Turing-based) PAC; non-CT cases are treated separately in the substrate section §8.
Data-path edits and i.i.d. integrity.
Our PAC statements assume $S$ and $V$ are i.i.d. from $\mathcal{D}$ and independent of each other. Permitted data-path operations are those that preserve i.i.d. draws (e.g., additional i.i.d. samples, balanced but label-independent subsampling, or predeclared splits). If selection depends on labels or on $V$, standard importance-weighting or covariate-shift corrections must be used; otherwise guarantees may fail.
Axes.
@@
    Algorithmic. Update rules, schedules, stopping, and internal randomness; the hypothesis family is fixed.
@@
    Representational. Changes to the hypothesis class or encoding, such as feature maps, basis expansions, or unions and refinements.
@@
The five-axis decomposition provides a conceptual toolkit for reasoning about arbitrary self-modification. Any concrete self-improving system can be analyzed by identifying which axes it modifies and tracing the induced changes to the hypothesis family. This has three immediate benefits: (i) it unifies seemingly disparate mechanisms under a common analytical framework; (ii) it isolates which modifications actually affect learnability (those that change the reachable hypothesis family) from those that affect only computational efficiency; and (iii) it enables modular safety certification—one can verify capacity bounds axis-by-axis rather than re-analyzing entire systems from scratch.
Concretely, these five axes already cover modern practice. Algorithmic edits appear in online hyperparameter and optimizer rewrites (Jaderberg et al.,, 2017). Representational expansions include mixture of experts and parameter-efficient adapters (Fedus et al.,, 2022; Houlsby et al.,, 2019; Hu et al.,, 2021). Architectural rewiring is explored by neural architecture search (Liu et al.,, 2019; Elsken et al.,, 2019). External memory, retrieval, and tool use expand effective computation while keeping computability assumptions unchanged (Graves et al.,, 2014, 2016; Lewis et al.,, 2020; Schick et al.,, 2023). Reflective and self-feedback agents instantiate metacognitive scheduling (Shinn et al.,, 2023; Madaan et al.,, 2023; Yao et al.,, 2023; Wang et al.,, 2023). As self-modifying capabilities become standard in deployed systems, this framework offers a principled way to predict, diagnose, and control when self-improvement preserves or destroys generalization guarantees.
@@
  1. 1.
    (A1) Data $(x,y)$ i.i.d. from fixed $\mathcal{D}$; training $S\sim P^{m}$ and validation $V\sim P^{n_{v}}$ are independent.
  2. 2.
    (A2) Loss $\ell\in[0,1]$.
  3. 3.
    (A3) Capacity is any uniform-convergence notion (VC/pseudodim/VC-subgraph); we instantiate VC where convenient.
  4. 4.
    (A4) When a computable proxy $B$ is used, it upper-bounds capacity: $B(\cdot)\geq\mathrm{cap}(\cdot)$.
  5. 5.
    (A5) Substrate semantics. If the substrate $F$ is Church–Turing equivalent, solvability/learnability are measured in the classical (Turing) sense; non-CT substrates (e.g., finite memory or oracular/analog) may alter this and are treated separately in §8.
  6. 6.
    (A6) Axis isolation. We analyze one axis at a time while holding the others fixed; multi-axis edits are discussed later.
  7. 7.
    (A7) Compute scope. We study sample complexity (learnability), not runtime, unless limits are intrinsic to $F$.
4 Representational Self-Modification ($\mathcal{M}_{H}$)
Setting (fixed vs. modifiable).

### v2

# On The Statistical Limits of Self-Improving Agents

@@ Abstract
As systems trend toward superintelligence, a natural modeling premise is that agents can self-improve along every facet of their own design. We formalize this with a five-axis decomposition and a decision layer, separating incentives from learning behavior and analyzing axes in isolation. Our central result identifies and introduces a sharp utility–learning tension, the structural conflict in self-modifying systems whereby utility-driven changes that improve immediate or expected performance can also erode the statistical preconditions for reliable learning and generalization. Our findings show that distribution-free guarantees are preserved iff the policy-reachable model family is uniformly capacity-bounded; when capacity can grow without limit, utility-rational self-changes can render learnable tasks unlearnable. Under standard assumptions common in practice, these axes reduce to the same capacity criterion, yielding a single boundary for safe self-modification.
@@ 1 Introduction
Evidence for this shift already exists. Reinforcement learning and meta-learning instantiate constrained self-change Sutton and Barto, 2018; Finn et al., 2017; Rajeswaran et al., 2019; Hospedales et al., 2022, while open-ended pipelines iterate code edits and tools Zhang et al., 2025. Decision-theoretic proposals investigate provably utility-improving modifications Schmidhuber, 2005, safety analyses document pathologies Orseau and Ring, 2011, and metagoal frameworks aim to stabilize goal evolution Goertzel, 2024. What remains underdeveloped is a learning-theoretic account of post-modification behavior: when do seemingly rational self-changes preserve the conditions under which learning is possible, and when do they destroy them? We prove a policy-level learnability boundary: distribution-free PAC guarantees are preserved if and only if the policy-reachable family has uniformly bounded capacity. A simple Two-Gate guardrail, validation margin $\tau$ plus capacity cap $K\!\left[m\right]$, keeps trajectories on the safe side and yields a VC-rate oracle inequality.
@@
  • Policy boundary with if and only if. Under standard i.i.d. assumptions, distribution-free PAC learnability is preserved under self-modification if and only if the policy-reachable family has uniformly bounded capacity, measured by VC dimension or an equivalent uniform-convergence notion.
  • Axis reductions. Architectural and metacognitive edits reduce to induced hypothesis families, and substrate changes matter only via computability or the induced family. Hence the boundary depends solely on the supremum capacity of the reachable family.
  • Two-Gate guardrail. A computable accept or reject rule, validation improvement by margin $\tau$ plus a capacity cap $K\!\left[m\right]$, ensures monotone true-risk progress and an oracle inequality at standard VC rates for the terminal predictor.
  • Axis clarity. We separate axes that expand the effective hypothesis family from axes that analyze structured subsets within standard models. In particular, the algorithmic axis is a structured subset within standard online learning, not a new model class.
@@ 2.1 Decision-theoretic self-modification and safety
Gödel Machines give a proof-based framework for globally optimal self-modification under a utility function Schmidhuber, 2005. Safety analyses document pathologies for self-modifying agents, including reward hacking and self-termination Orseau and Ring, 2011. Open-ended empirical systems iterate code and toolchain edits with benchmark gains but without proof obligations Zhang et al., 2025. Proposals for metagoals aim to stabilize or moderate goal evolution during self-change Goertzel, 2024. While these frameworks establish decision-theoretic foundations, they do not provide learning-theoretic guarantees about post-modification generalization. We study the learning-theoretic state of the agent after such edits.
@@ 2.2 Modern mechanisms for self-improvement
Contemporary machine learning exhibits constrained forms of self-modification across multiple dimensions. Neural architecture search explores architectural topologies through differentiable, evolutionary, and reinforcement approaches Liu et al., 2019; Elsken et al., 2019; Zoph and Le, 2017; Real et al., 2019. Automated machine learning systems perform pipeline and hyperparameter search and can trigger optimizer and model-family switches Hutter et al., 2019; Feurer and Hutter, 2019; Li et al., 2017. Population-based training simultaneously evolves hyperparameters and weights across a population of models Jaderberg et al., 2017.
Meta-learning adapts optimizers, initializations, and inductive biases across tasks Finn et al., 2017; Rajeswaran et al., 2019; Hospedales et al., 2022. Reinforcement learning and multi-armed bandits provide policies for selecting modifications and exploration strategies Sutton and Barto, 2018; Auer et al., 2002; Lai and Robbins, 1985; Slivkins, 2019. Representation growth through mixture of experts and adapters, and the use of external memory and retrieval, expand the effective function family and computation available at inference Fedus et al., 2022; Houlsby et al., 2019; Hu et al., 2021; Graves et al., 2014, 2016; Lewis et al., 2020; Schick et al., 2023. Continual learning addresses sequential task acquisition while mitigating catastrophic forgetting Kirkpatrick et al., 2017; Parisi et al., 2019; Van de Ven and Tolias, 2022.
These mechanisms instantiate partial self-modification, adapting specific components while keeping the learning framework itself fixed. In contrast, true self-modifying agents can rewrite any axis of their design. In our framework, these mechanisms traverse representational, architectural, algorithmic, and metacognitive axes, and we establish when such traversals preserve or destroy learnability.
@@ 2.3 Learning theory for adaptive systems
PAC learning provides distribution-free guarantees under a fixed hypothesis class and algorithm Shalev-Shwartz and Ben-David, 2014; Mohri et al., 2018; Blumer et al., 1989; Vapnik, 1998; Hanneke et al., 2024. Online learning theory establishes regret bounds for adaptive algorithms Shalev-Shwartz, 2012; Hazan, 2016; Cesa-Bianchi and Lugosi, 2006, but assumes the learning mechanism remains fixed. Transformation-invariant learners extend instance equivalence while keeping the learning mechanism fixed Shao et al., 2022. Predictive PAC relaxes data assumptions with a fixed learner Pestov, 2010, and iterative improvement within constrained design spaces admits PAC-style analysis Attias et al., 2025.
Information-theoretic approaches bound generalization for adaptive and meta-learners via mutual information Jose and Simeone, 2021; Chen et al., 2021; Wen et al., 2025. Stability connects optimization choices to generalization bounds Bousquet and Elisseeff, 2002; Hardt et al., 2016. All of these results assume architectural invariance: the hypothesis class, update rule, or computational model is fixed ex ante. We remove this assumption and characterize when self-modification preserves PAC learnability.
@@ 2.4 Computability and the substrate
Church–Turing equivalent substrates preserve solvability up to simulation overhead, whereas strictly weaker substrates with finite memory can forfeit learnability of classes that are otherwise PAC-learnable. Stronger-than-Turing models change the problem class under discussion Akbari and Harrison-Trainor, 2024. This motivates treating substrate edits separately from architectural or representational changes and clarifies when invariance should be expected.
@@ 3 Setup and Five-Axis Decomposition
Data and splits.
We study supervised learning under an unknown distribution $\mathcal{D}$ over examples $\langle x,y\rangle\in\mathcal{X}\times\mathcal{Y}$. Let $S\sim\mathcal{D}^{m}$ denote the training sample and $V\sim\mathcal{D}^{n_{v}}$ an independent validation sample.
Loss, risk, and utility are distinct objects.
Fix a loss $\ell:\mathcal{Y}\times\mathcal{Y}\to\left[0,1\right]$ and define population risk $R\!\left[h\right]=\mathbb{E}_{\langle x,y\rangle\sim\mathcal{D}}\!\left[\ell\!\left[h\!\left[x\right],y\right]\right]$ and empirical risks $\widehat{R}_{S}$ and $\widehat{R}_{V}$. Utility $u$ is an agent-internal objective used to decide whether to self-modify. It may depend on proxies of risk, resources, or constraints. We do not assume $u$ equals negative loss unless stated.
Learner state and axes.
@@
$\ell_{t}=\langle A_{t},H_{t},Z_{t},F_{t},M_{t}\rangle\in\underbrace{\mathcal{L}}_{\mathcal{A}\times\mathcal{H}\times\mathcal{Z}\times\mathcal{F}\times\mathcal{M}},$
where $A$ is algorithmic, $H$ is representational, $Z$ is architectural, $F$ is substrate, and $M$ is metacognitive.
Finite evidence and the modification map.
Let $\mathcal{E}$ denote the space of finite evidence objects available to the agent, such as a current minibatch, a fixed buffer of past data, a predeclared split of $S$, or other finite statistics derived from the interaction history. At time $t$ the agent has finite evidence $E_{t}\in\mathcal{E}$. A possibly stochastic modification map $\Phi:\mathcal{L}\times\mathcal{E}\to\mathcal{L}$ updates the system via
$\ell_{t+1}=\Phi\!\left[\ell_{t},E_{t}\right].$
For $X\in\left\{A,H,Z,F,M\right\}$ with state space $\mathcal{X}$, we write
$X_{t+1}=\Phi_{X}\!\left[X_{t},E_{t},\theta_{X,t}\right],\qquad\theta_{X,t}\in\Theta_{X},$
where $\Theta_{X}$ indexes admissible edits along axis $X$.
Environment context.
We allow utility to depend on an external environment context $\mathrm{Env}_{t}$, such as available compute, wall-clock time, or deployment constraints, in addition to finite evidence and internal state.
Decision rule and tie-breaking.
A candidate modification at time $t$ is executed if and only if there exists a formal proof in the agent’s current calculus that it yields an immediate utility increase:
$u\!\left[\Phi\!\left[\ell_{t},E_{t}\right],\mathrm{Env}_{t}\right]\;>\;u\!\left[\ell_{t},\mathrm{Env}_{t}\right].$
If multiple candidates satisfy this criterion, selection is handled by the metacognitive scheduler $M_{t}$, such as first-found proof, maximum certified utility gain, or any fixed tie-break rule. Our theorems are stated in terms of the resulting policy-reachable set induced by this selection rule.
Policy-reachable families.
For any axis $X\in\left\{A,H,Z,F,M\right\}$, let $\mathcal{X}_{\mathrm{reach}}\!\left[u\right]$ be the set of $X$ states appearing along some trajectory generated by the decision semantics from $\ell_{0}$ under utility $u$.
Learnability, distribution-free PAC and computable PAC.
A hypothesis family $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ is distribution-free PAC learnable if there exists a learner $\mathsf{Alg}$ such that for all $\varepsilon,\delta\in\left]0,1\right[$, for all distributions $\mathcal{D}$, and for $m\geq m_{\mathcal{H}}\!\left[\varepsilon,\delta\right]$, if $S\sim\mathcal{D}^{m}$ then with probability at least $1-\delta$ the output $\hat{h}=\mathsf{Alg}\!\left[S\right]$ satisfies
$R\!\left[\hat{h}\right]\leq\inf_{h\in\mathcal{H}}R\!\left[h\right]+\varepsilon.$
When we say learnable in the Turing sense, we additionally require that the learner and the hypotheses it outputs are implementable on a Church–Turing equivalent substrate.
@@
Our results apply with any uniform capacity notion that yields distribution-free uniform convergence, such as VC-subgraph or pseudodimension. For concreteness, for 011 loss we use VC dimension, written $\mathrm{VC}\!\left[\cdot\right]$.
@@
We use absolute constants $c_{1},c_{2},\ldots$ whose values may differ across lemmas but are fixed within a statement. We use $\tilde{O}\!\left$$\cdot\right$] to hide polylogarithmic factors in $m,n_{v},1/\delta$. Probabilities are over the draws of $S$ and $V$ unless specified.
@@
Throughout, we analyze one axis at a time while holding others fixed. Under Church–Turing equivalent substrates $F$, learnability refers to classical PAC in the computable sense. Non Church–Turing cases are treated separately in Section 8.
Data-path integrity.
Our PAC statements assume $S$ and $V$ are i.i.d. from $\mathcal{D}$ and independent of each other. Permitted data-path operations are those that preserve i.i.d. draws, such as additional i.i.d. samples, balanced but label-independent subsampling, or predeclared splits. If selection depends on labels or on $V$, standard importance-weighting or covariate-shift corrections must be used; otherwise guarantees may fail.
Axes and interpretation.
Each axis corresponds either to expanding the effective hypothesis family or to analyzing a structured subset within a standard model. Representational and architectural axes expand the effective hypothesis family. The algorithmic axis studies a structured subset within standard online learning. The metacognitive axis filters which edits are evaluated and accepted. The substrate axis affects learnability only through computability or by changing the induced hypothesis family.
@@
    Algorithmic. Update rules, schedules, stopping, and internal randomness. This axis is a structured subset within standard online learning.
@@
    Representational. Changes to the hypothesis class or encoding, such as feature maps, basis expansions, unions, and refinements.
@@
Self-improvement is often discussed as a monolith, but it is not: agents can change what they can represent, how they search, how information flows, what compute they have, and how they choose among modifications. The five-axis decomposition makes this explicit and, crucially, makes it analyzable. Each axis induces a set of post-modification predictors, and the union of what the agent can reach under its decision rule is the only object that matters for distribution-free guarantees. Once phrased this way, the safety question becomes concrete: does self-improvement keep the reachable family capacity-bounded, or does it allow capacity to drift without limit? The decomposition therefore turns an amorphous concern—agents that rewrite themselves—into a small number of verifiable invariants.
@@
  1. A1
    Data $\langle x,y\rangle$ are i.i.d. from fixed $\mathcal{D}$. Training $S\sim\mathcal{D}^{m}$ and validation $V\sim\mathcal{D}^{n_{v}}$ are independent.
  2. A2
    Loss is bounded, $\ell\in\left[0,1\right]$.
  3. A3
    Capacity is any uniform-convergence notion, such as VC, pseudodimension, or VC-subgraph. We instantiate VC where convenient.
  4. A4
    When a computable proxy $B$ is used, it upper-bounds capacity, $B\!\left[\cdot\right]\geq\mathrm{cap}\!\left[\cdot\right]$.
  5. A5
    Substrate semantics. If the substrate $F$ is Church–Turing equivalent, solvability and learnability are measured in the classical computable sense. Non Church–Turing substrates may alter this and are treated separately in Section 8.
  6. A6
    Axis isolation. We analyze one axis at a time while holding the others fixed. Multi-axis edits are discussed later.
  7. A7
    Compute scope. We study sample complexity, not runtime, unless limits are intrinsic to $F$.
4 Representational Self-Modification $\mathcal{M}_{H}$
Setting, fixed versus modifiable.